a disc of mass ten gram is kept horizontally in air by firing bullets of mass five g each at rate of ten per second.if the bullets rebound with same speed .The velicity with which bullets are fired is

Grade:12

1 Answers

Arun
6 years ago
Dear student
 
force is 
F = change of momentum/ time = impulse/time 
the change of momentum of the bullet is 
impulse = vm - (- v)m = 2vm = 2v*0.005 kgm/s 
and assuming that the time of impact of 1 bullet is 0.1 second, then the force is 
F = 2v*0.005/0.1 N 
now set weight = Force to find v: 
0.01 kg * 9.81 m/s^2 = ( 2*0.0005 kg*/ 0.1 s) *v
v = 0.01*9.81*0.1/(2*0.005) 
v = 0.98 m/s (= 98 cm/s)

Regards
Arun (askIITians forum expert)